Frequently Asked Questions
- Special Thunderbird IMAP Settings
Special IMAP settings in Thunderbird help to prevent corrupted attachments resulting from the behavior of certain IMAP servers such as Gmail and instances of MS Exchange Server which have a particular 'shortcut' feature turned on.
- New Login Accounts
How to request a new login account.
- How to set up the lecture room video projectors
This contains a link to an information sheet with steps and diagrams explaining the setup of the video projectors in the large BPS Bldg. lecture halls.
- Finding information on P-A Department Printers
- Changing P-A main web server passwords
sshto log into the
- Change Your Password in Outlook Web App / Access
How to change your PHY-AST domain and Email password on Outlook Web App / Outlook Web Access.
|Campus Internet/Network Issues|
|Article # 81 - Posted by Ehren Benson on 2013-09-25 14:54:31|
|Around 1pm this afternoon we started to notice network issues in the building, we called it in and were told by ITS that there are issues all over campus, at this point everything INSIDE the building is working but it is not possible to get out, or back in from the outside.|
|Web services downtime|
|Article # 80 - Posted by George Perkins on 2013-08-23 18:53:15|
On Monday, 26 August 2013, starting at about 10:10 am, many of the P-A Departmental web sites will be down for a short time (we expect it to be less than 30 minutes) to replace a failing component in one of the servers. These will include the main P-A web site, this support site, the Science Theatre site, and several others.
|Exchange Server Outages|
|Article # 79 - Posted by Ehren Benson on 2013-08-08 16:27:51|
|One of the 3 exchange server database servers will be down for the evening to expand the database storage volumes which takes some time. There are two other available servers to handle requests but you may occasionally have trouble connecting. If so give it a little time and try again. Thanks.|
|Server room network problem this evening|
|Article # 78 - Posted by George Perkins on 2013-06-18 20:39:40|
This evening (Tuesday, 18 June 2013) around 5 pm, one of the two main network switches in the P-A server room developed a serious problem and locked out network traffic from roughly half of the servers in the room. We had it back up again around 6:20 pm, but it became clear afterwards that it was failing in more than one way, so network connections were moved over to alternate switches to make sure that departmental services were available, even if potentially on slower-than-usual network links temporarily.
Tomorrow morning, some of the systems now on a lower speed temporary switch will be moved to a higher speed temporary switch, and the manufacturer of the faulty switch will be contacted for repair or replacement (probably the latter).
Services affected included many of the department's web sites and its incoming E-mail server. Some outgoing E-mail may have been also blocked during the network outage, as that feature is shared between three different servers, one of which was connected to the outside world via the bad switch while the other two were not. One may have experienced delays in reading E-mail on the server, but that service should not have failed entirely (though particular connections may have been dropped and been forced to be re-established; some E-mail reading software handles this better than others).